The sum of the composite numbers between 20 and 30,minus the sum of the sum of the prime numbers
Answer by LinnW(1048)

Composite numbers between 20 and 30 are
20 21 22 24 25 26 27 28 30
There sum is 20+21+22+24+25+26+27+28+30 = 223
Prime numbers between 20 and 30
23 29
Their sum is 23+29 = 52
223 - 52 = 171